A negative number decreased by its reciprocal is -24/5. Find the number.
Answer by Cromlix(4381)

Negative number = -x
A negative number decreased by its reciprocal is -24/5.
-x - (-1/x) = -24/5
-x + 1/x = -24/5
Common denominator = 5x
-5x^2 + 5 = - 24x
Collect terms;
-5x^2 + 24x + 5 = 0
(-5x - 1)(x - 5) = 0
...........
-5x - 1 = 0
-5x = 1
x = -1/5 (no answer)
............
x - 5 = 0
x = 5
-5 - (-1/5) = -24/5
Hope this helps.
:-)